OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, give the limit. 48, 12, 3, 3/4, ...

OpenStudy (anonymous):

Is it increasing or decreasing?

OpenStudy (anonymous):

converges if it looks like it will approach a certain number (limit)

OpenStudy (anonymous):

Converges; 64 Diverges Converges; -4080 Converges; 0

OpenStudy (anonymous):

Answer it . Not show me the choices

OpenStudy (anonymous):

decreasing

OpenStudy (anonymous):

Do you know by what pattern?

OpenStudy (anonymous):

its dividing by 4

OpenStudy (anonymous):

Yea you are right. After 3/4 it comes to \[\frac{ 3 }{ 16 }, \frac{ 3 }{ 64 }, \frac{ 3 }{ 256 }\] ............ since 3/256 or 0.117 appears to be a very small number and the pattern continues to be smaller, it approaches zero Therefore it converges at zero

OpenStudy (anonymous):

oh ok i understand.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

No problem!